OpenStudy (anonymous):

Can anyone tell me how to do this? Just walk me through it: Find all self-crossings if any for the following curve: x(t) = 1+t^2, y(t) = t^3

OpenStudy (anonymous):

A point of self crossing means for more than one value of t yields the same (x,y)

OpenStudy (anonymous):

Ok, so how should I begin to find that?

OpenStudy (anonymous):

Let t1 and t2 be values of t such that the same values of x and y result: \[x(t _{1})=x(t _{2}) ; y(t _{1})=y(t _{2})\] Therefore:\[1+t _{1}^{2}=1+t _{2}^{^{2}} ; t _{1}^{^{3}}=t _{2}^{3}\]

OpenStudy (anonymous):

So we need to determine if it is possible to have distinct values t1 and t2 which satisfy the above equations. Right?

OpenStudy (anonymous):

Got it.

OpenStudy (anonymous):

And then how do I proceed?

OpenStudy (anonymous):

As can be seen, if \[t _{1}^{3}=t _{2}^{3}\] then it must be that t1=t2. Right?

OpenStudy (anonymous):

We cannot have distinct values of t1and t2. So to me it seems that this curve does not self intersect.

OpenStudy (anonymous):

Do you have the answer? To me it seems this curve does not cross itself and a quick check using graphing tools seems to show the same. Let me know if you believe there is an error in this as per your information. Thanks.

OpenStudy (anonymous):

It seems to be so. Thank you.